# Overlap Gap and Computational Thresholds in the Square Wave Perceptron

### v2

@@ A.2 The Truncated Square Wave
$\begin{split}I_{K,\gamma}(m,\sigma)&\equiv\int Dh\,\Theta\left(\varphi(\sigma\,h+m)\right)=\int_{\frac{\gamma-m}{\sigma}}^{\infty}Dh+\sum_{l=0}^{K-1}\int_{\frac{-\gamma+\frac{2l\gamma}{K}-m}{\sigma}}^{\frac{-\gamma+\frac{(2l+1)\gamma}{K}-m}{\sigma}}Dh=\\ &=H\left(\frac{\gamma-m}{\sigma}\right)+\sum_{l=0}^{K-1}\left(H\left(\frac{-\gamma+\frac{2l\gamma}{K}-m}{\sigma}\right)-H\left(\frac{-\gamma+\frac{(2l+1)\gamma}{K}-m}{\sigma}\right)\right)\,,\end{split}$ (91)
@@ A.3 The Square Wave
[…]
@@
Appendix B The Replica-Symmetric Free Entropy of One Clone
@@
B.1 Annealed Upper Bound on the Storage Capacity
@@
B.2 Annealed Upper Bound on the Teacher-Student
@@
Appendix C Satisfiability threshold in the small $\delta$ limit
@@
[…]
Appendix D Number of $m$-Cliques in the Large $m$ Limit
@@
Appendix E Expansion for small $\delta$
@@
In Fig. 11(a) we compare the small $\delta$ expansion of the annealed OGP thresholds $\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)\approx-G^{(m)}_{S}(\varepsilon)/G^{(m)}_{E}(\varepsilon)$, with the estimates $\alpha^{\text{ann}}_{OGP}(m,\delta)$ obtained from the numerical solution of (11). Note that $\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)\geq\alpha^{\text{ann}}_{OGP}(m,\delta)$, therefore we can use $\min_{m}\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)$ as an upper bound to the OGP threshold. In Fig. 11(b) we represent the curve $\min_{m}\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)$.
@@
Appendix F Estimating the $m$-OGP using a Replica-Symmetric Ansatz
@@
The expert reader may recognize that imposing an RS ansatz on the clonated free entropy (8) is effectively equivalent to imposing a 1-step Replica Symmetry Breaking ansatz (1RSB) on the equilibrium free entropy. The key difference, however, is that in computing (8), one must fix both the Parisi block size $m$ and the overlap $q_{1}$, rather than optimizing over them as in standard equilibrium calculation. Moreover, $m$ need not to be in the physical range $10<m<$ 1. In the limit $n\to 0$, one gets,
@@
[…]
@@
$\phi^{\star}(m)$ $=-\frac{\partial(m\phi_{m})}{\partial m}\,,$ (143a)
@@
If $\Sigma(m)<0$ for some $m$ this signals that the solution we found is nonphysical. This usually happens if $m$ is large enough. We show an example of this in Fig. 12, in the case of the square wave activation with $\delta=1.5$.
@@
Appendix G Reinforced Approximate Message Passing
@@
[…]
@@
We call “iteration” an update of all magnetizations (and therefore of all the latent fields). The parameter $r$ in (149) is called reinforcement rate. Setting $r=0$, RAMP becomes equivalent to the approximate message passing (AMP) algorithm, which is used to estimate the local marginals of variables in constraint-satisfaction problems [21]. In the teacher-student case, for $\alpha\geq\alpha_{r}$, the local magnetizations polarize in the direction of the teacher. An example is given in Figs. 8(a) and 8(b). This is the only situation where AMP provides a solver for Eq. 144. In other cases, AMP returns non-integer magnetization that do not define a vertex of the hypercube. Setting $r>0$, one introduces in the AMP iteration a term that tends to polarize the magnetizations, providing a heuristic solver for the problem [64, 48]. In this case, equations (145)-(150) are iterated until the binary vector $\tilde{\boldsymbol{x}}^{t}$, with elements $\tilde{x}^{t}_{i}=\text{sgn}(m_{i}^{t})$, satisfies constraint 144, or $t$ becomes larger that a maximum time $T_{max}$, that is a parameter of the algorithm. There are different choices of the reinforcement term suggested in the literature, which all lead to qualitatively similar results. The behavior of the solver depends on the choice of $r$. The performance of the solver, which is determined by the largest value of $\alpha$ at which the solver is able to find solutions in polynomial time, improves by reducing $r$. However, it is also found that the number of iterations required to find a solution scales like $1/r$. Therefore, it is convenient to define the optimum reinforcement rate $r_{opt}$, which is the largest (therefore the cheapest) reinforcement rate that allows finding a solution [48, 26]. In Fig. 5 we show $T_{sol}$, which is the average number of iterations required to find a solution when the reinforcement rate is chosen as the optimal one, as a function of $\alpha$ for different values of $\delta$.

### v4

@@ A.2 The Truncated Square Wave
[…]
@@ A.3 The Square Wave
$b_{n}=\frac{2}{T}\int_{0}^{T}dx\,\varphi(x)\,\sin{\frac{2\pi nx}{T}}=\frac{1}{\delta}\int_{\delta}^{2\delta}dx\,\sin{\frac{\pi nx}{\delta}}-\frac{1}{\delta}\int_{0}^{\delta}dx\,\sin{\frac{\pi nx}{\delta}}=\frac{4\cos{(n\pi)}\,\sin^{2}{(\frac{n\pi}{2}})}{n\pi},\quad n\geq 1\,.$ (97)
@@
Appendix B Cloning in different instances of disorder
In a variation of the $m$-OGP, called ensemble OGP (e-OGP), the clones are allowed to live in different instances of the disorder. This variation of OGP is used to construct a refutation to stable algorithms.
Consider for simplicity the storage case with $m=2$. We denote by $\boldsymbol{\xi}^{\mu,1}$ and $\boldsymbol{\xi}^{\mu,2}$ the realization of pattern $\mu$ in the first and second instance, respectively. We consider:
$\mathds{E}\,\xi_{i}^{\mu,1}=\mathds{E}\,\tilde{\xi}_{i}^{\mu,2}=0,\quad\mathds{E}\,\big(\xi_{i}^{\mu,1}\big)^{2}=\mathds{E}\,\big(\tilde{\xi}_{i}^{\mu,2}\big)^{2}=1,\quad\mathds{E}\,\xi_{i}^{\mu,1}\tilde{\xi}_{j}^{\nu,2}=\delta_{ij}\delta_{\mu\nu}\rho,$ (101)
for some $\rho\in[0,1]$, where $\mathds{E}$ is the expectation over the patterns. The energetic term of the clonated free entropy for $m=2$ reads:
$G_{E}^{(2)}=\frac{1}{2}\,\log{\int D_{\Sigma}\boldsymbol{u}}\,\Theta(\varphi(u_{1}))\,\,\Theta(\varphi(u_{2})),$ (102)
where the elements of the covariance matrix $\Sigma$ are given by:
$\Sigma_{ab}=\delta_{ab}+\rho\,q\,(1-\delta_{ab}).$ (103)
Note that the single-instance case that we analyzed in the paper is obtained taking $\rho\to 1$. We call $I(\rho\,q)=\exp{(2\,G_{E}^{(2)})}$, and $r=\rho\,q$. Let $s(z)=\mathrm{sign}(\varphi(z))\in\{\pm 1\}$ denote the zero-mean square-wave sign so that $\Theta(\varphi(z))=\tfrac{1}{2}(1+s(z))$. We define
$I(r)=\mathbb{E}\!\left[\Theta(\varphi(u_{1}))\,\Theta(\varphi(u_{2}))\right]=\frac{1}{4}\Big(1+C(r)\Big),\qquad C(r):=\mathbb{E}[s(u_{1})\,s(u_{2})].$
Every square-integrable function $f$ of a standard Gaussian variable admits an expansion in the orthonormal Hermite basis,
$f(z)=\sum_{n\geq 0}a_{n}\,h_{n}(z),\qquad h_{n}(z):=\frac{H_{n}(z)}{\sqrt{n!}},\qquad a_{n}=\mathbb{E}\big[f(Z)\,h_{n}(Z)\big],$
where $H_{n}(z)$ are the (probabilists’) Hermite polynomials and $Z\sim\mathcal{N}(0,1)$. For the odd function $s(z)$, all even coefficients vanish, $a_{2k}=0$. Writing $s(u_{a})=\sum_{n\geq 0}a_{n}\,h_{n}(u_{a})$ for $a\in\{1,2\}$ and using the standard identity for correlated Gaussians
$\mathbb{E}\!\left[h_{n}(u_{1})\,h_{m}(u_{2})\right]=\delta_{nm}\,r^{\,n},$
we obtain
$\begin{split}C(r)&=\mathbb{E}[s(u_{1})s(u_{2})]=\sum_{n,m}a_{n}a_{m}\,\mathbb{E}[h_{n}(u_{1})h_{m}(u_{2})]=\sum_{n\geq 0}a_{n}^{2}\,r^{\,n}=\sum_{k\geq 0}a_{2k+1}^{2}\,r^{\,2k+1}.\end{split}$
Differentiating with respect to $r$ gives
$C^{\prime}(r)=\sum_{k\geq 0}(2k{+}1)\,a_{2k+1}^{2}\,r^{\,2k}\geq 0,$
with strict inequality for nontrivial $s$. This means that $G_{E}^{(2)}(\rho\,q)$ is a strictly increasing function of $\rho\in[0,1]$, implying that
$\alpha_{2OGP}^{\text{ann}}(\rho)\leq\alpha_{2OGP}^{\text{ann}}(1),$ (104)
where $\alpha_{2\mathrm{OGP}}^{\text{ann}}(\rho)$ denotes the annealed 2-OGP threshold computed with the two clones living in correlated instances, and $\alpha_{2\mathrm{OGP}}^{\text{ann}}(1)$ the corresponding threshold when both clones belong to the same instance. In other words, the presence of an OGP within the same instance implies its presence across two instances with correlation defined in Eq. (101). In this sense, the single-instance setting represents a worst-case scenario, which justifies our focus on this case. A similar argument can be done for $m>2$ assuming uniform correlation between the different $m$ instances.
Appendix C The Replica-Symmetric Free Entropy of One Clone
@@
C.1 Annealed Upper Bound on the Storage Capacity
@@
C.2 Annealed Upper Bound on the Teacher-Student
@@
Appendix D Satisfiability threshold in the small $\delta$ limit
@@
Figure 9: (a): Second moment free entropy $\max_{q}\phi^{\text{ann}}_{2}(q,\alpha)$ as a function of $\alpha$ for different values of $\delta$. From bottom to top $\delta=1.2+n\,0.2$, with $n=0,\dots,19$. The dashed line corresponds to $(1-\alpha)\log{2}$. For $\delta\rightarrow 0$, $(1-\alpha)\log{2}-\max_{q}\phi^{\text{ann}}_{2}(q,\alpha)\rightarrow 0$ for all values of $0<\alpha<1$. Inset: $\max_{q}\phi^{\text{ann}}_{2}(q,\alpha=1)$ as a function of $\delta$. (b): optimal overlap of the second moment free entropy, $\text{argmax}\phi_{1}^{ann}(q,\alpha)$, as a function of $\alpha$ for different values of $\delta$. From bottom to top $\delta=1.2+n\,0.2$, with $n=0,\dots,19$. Note that reducing $\delta$, also the optimal overlap reduces for all values of $\alpha$.
Appendix E Non-monotonicity of the hard–easy transition
While the divergence of the hard–easy transition as $\delta\to 0$ admits a simple interpretation in terms of disorder observables (see Sec. 4.3.2), the initial decrease of the transition when lowering $\delta$ from $\infty$ is more subtle. This feature can be seen directly from the RS formulas by comparing the energetic terms of the free entropies. Indeed, we can decompose the RS potential as
$\Phi_{\text{SWP}}(p,\alpha)\;=\;G_{S}(p)\;+\;\alpha\,G_{E}^{\text{SWP}}(p)\;=\;\Phi_{\text{ABP}}(p,\alpha)\;+\;\alpha\,\Delta G_{E}(p,\delta),$ (123)
where $\Phi_{\text{SWP}}$ and $\Phi_{\text{ABP}}$ are the free entropies of the square–wave and asymmetric binary perceptron respectively, on the Nishimori line, and $\Delta G_{E}(p,\delta)=G_{E}^{\text{SWP}}(p,\delta)-G_{E}^{\text{ABP}}(p)$, where $G_{E}^{\text{SWP}}(p,\delta)$ and $G_{E}^{\text{ABP}}(p)$ are obtained from equation A.26 by substituting respectively $\varphi(h)=-\mathrm{sign}\left(\sin\left(\frac{\pi h}{\delta}\right)\right)$ and $\varphi(h)=\mathrm{sign}(h)$, and taking $p=q$ (Nishimori condition).
Figure 10: From top left to bottom left: $\alpha\Delta G_{E}$, $\Phi_{ABP}$ and $\Phi_{SWP}$ computed for $\alpha=\alpha_{ABP}$, $\delta=2.1$. Note that $\Phi_{SWP}$ is monotonously increasing, meaning that AMP messages initialized in a completely uninformed way ($p=0$) would converge to the teacher. This means that $\alpha_{ABP}$ is in the easy phase for SWP with $\delta=2.1$.
For $\alpha=\alpha_{\text{ABP}}$ (the ABP hard–easy transition point), one has $d\Phi_{\text{ABP}}/dp>0$ for all overlaps $p\neq p_{\text{ABP}}$, and $d\Phi_{\text{ABP}}/dp=0$ only at the ABP maximizer $p_{\text{ABP}}\simeq 0.89$ (with $\alpha_{\text{ABP}}\simeq 1.49$, see [71]). By analyzing the correction term $\Delta G_{E}$, one finds that for sufficiently large $\delta$ the total derivative
$\frac{d\Phi_{\text{SWP}}}{dp}=\frac{d\Phi_{\text{ABP}}}{dp}+\alpha^{\text{ABP}}\,\frac{d\Delta G_{E}}{dp}$
remains strictly positive for all values of the overlap $p$. In particular, this implies $d\Delta G_{E}(p^{\text{ABP}},\delta)/dp>0$. The global positivity excludes the appearance of spurious stationary points and guarantees that the SWP hard–easy transition must occur at a strictly smaller value of $\alpha$ than in ABP.
This reasoning is illustrated in Fig. 10, where we plot $\Phi_{\text{SWP}}(p)$, $\Phi_{\text{ABP}}(p)$, and the correction term $\alpha\Delta G_{E}(p,\delta)$ at $\alpha=\alpha_{\text{ABP}}$ and $\delta=2.1$. We stress that $\Delta G_{E}(p,\delta)$ is a simple function of $p$ that can be studied without any further optimization. Providing a more geometric interpretation of the behavior of $\Delta G_{E}$ would be interesting, but we leave this point for future work.
We note that, in contrast to AMP, the information–theoretic threshold decreases monotonically with $\delta$ and approaches one as $\delta\to 0$, since finer square–wave constraints progressively eliminate other vertices of the hypercube that reproduce the same labels as the teacher at lower values of $\alpha$. In the limit $\delta\to 0$, the teacher becomes information–theoretically identifiable as soon as $\alpha>1$, for instance by exhaustive search. Thus, the best possible algorithm, in terms of the number of examples required to recover the teacher, improves its performance as $\delta$ decreases. A possible explanation of the non–monotonic behavior of AMP is that the algorithm can only exploit a restricted part of the available signal: for large $\delta$ most of the information lies in low modes that AMP can capture, while decreasing $\delta$ gradually shifts weight to higher modes. This may initially improve the effective performance of AMP, mimicking the optimal case, but eventually degrades it once the relevant information is concentrated in modes beyond its reach, thereby producing the observed non–monotonicity.
Appendix F Number of $m$-Cliques in the Large $m$ Limit
@@
Appendix G Expansion for small $\delta$
@@
In Fig.˜12(a) we compare the small $\delta$ expansion of the annealed OGP thresholds $\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)\approx-G^{(m)}_{S}(\varepsilon)/G^{(m)}_{E}(\varepsilon)$, with the estimates $\alpha^{\text{ann}}_{OGP}(m,\delta)$ obtained from the numerical solution of (11). Note that $\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)\geq\alpha^{\text{ann}}_{OGP}(m,\delta)$, therefore we can use $\min_{m}\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)$ as an upper bound to the OGP threshold. In Fig.˜12(b) we represent the curve $\min_{m}\widetilde{\alpha}^{\text{ann}}_{OGP}(m,\delta)$.
@@
Appendix H Estimating the $m$-OGP using a Replica-Symmetric Ansatz
@@
The expert reader may recognize that imposing an RS ansatz on the clonated free entropy (8) is effectively equivalent to imposing a 1-step Replica Symmetry Breaking ansatz (1RSB) on the equilibrium free entropy. The key difference, however, is that in computing (8), one must fix both the Parisi block size $m$ and the overlap $q_{1}$, rather than optimizing over them as in standard equilibrium calculation. Moreover, $m$ need not to be in the physical range $0<m<1$. In the limit $n\to 0$, one gets,
@@
$\mathcal{N}_{m}=\sum_{\alpha}\mathrm{e}^{Nm\phi_{\alpha}}=\int d\phi\sum_{\alpha}\delta(\phi-\phi_{\alpha})\mathrm{e}^{Nm\phi}=\int d\phi\,\mathrm{e}^{N(m\phi+\Sigma(\phi))}\,$, (147)
@@
$\phi^{\star}(m)$ $=\frac{\partial(m\phi_{m})}{\partial m}\,$, (150a)
@@
If $\Sigma(m)<0$ for some $m$ this signals that the solution we found is nonphysical. This usually happens if $m$ is large enough. We show an example of this in Fig.˜13, in the case of the square wave activation with $\delta=1.5$.
@@
Appendix I Reinforced Approximate Message Passing
@@
$f(\omega,V,y)=\partial_{\omega}\log{\int dz\,\mathrm{e}^{-\frac{(z-\omega)^{2}}{2V}}\Theta\big(y\,\varphi(z)\big)}\,.$ (158)
@@
We call “iteration” an update of all magnetizations (and therefore of all the latent fields). The parameter $r$ in (156) is called reinforcement rate. Setting $r=0$, RAMP becomes equivalent to the approximate message passing (AMP) algorithm, which is used to estimate the local marginals of variables in constraint-satisfaction problems [29]. In the teacher-student case, for $\alpha\geq\alpha_{r}$, the local magnetizations polarize in the direction of the teacher. An example is given in Figs.˜8(a) and 8(b). This is the only situation where AMP provides a solver for Eq.˜151. In other cases, AMP returns non-integer magnetization that do not define a vertex of the hypercube. Setting $r>0$, one introduces in the AMP iteration a term that tends to polarize the magnetizations, providing a heuristic solver for the problem [72, 34]. In this case, equations (152)-(157) are iterated until the binary vector $\tilde{\boldsymbol{x}}^{t}$, with elements $\tilde{x}^{t}_{i}=\text{sgn}(m_{i}^{t})$, satisfies constraint 151, or $t$ becomes larger that a maximum time $T_{max}$, that is a parameter of the algorithm. There are different choices of the reinforcement term suggested in the literature, which all lead to qualitatively similar results. The behavior of the solver depends on the choice of $r$. The performance of the solver, which is determined by the largest value of $\alpha$ at which the solver is able to find solutions in polynomial time, improves by reducing $r$. However, it is also found that the number of iterations required to find a solution scales like $1/r$. Therefore, it is convenient to define the optimum reinforcement rate $r_{opt}$, which is the largest (therefore the cheapest) reinforcement rate that allows finding a solution [34, 33]. In Fig.˜5 we show $T_{sol}$, which is the average number of iterations required to find a solution when the reinforcement rate is chosen as the optimal one, as a function of $\alpha$ for different values of $\delta$.